Susie has designed an exercise program for herselfS. One part of the program requires her to walk between 25 and 30 miles each week.She plans to walk the same distance each day five days a week. What is the range of miles that she should walk each day?
Answer by Theo(13342)

25/5 = 5
30/5 = 6

she should walk between 5 and 6 miles per day.

if 5 every day, then she walked 25.
if 6 every day, then she walked 30.

if anything between 5 and 6, then her total will be between 25 and 30.

for example - 5.5 * 5 = 27.5

the range of miles she should walke each day is 5 <= x <= 6, where x is the number of miles she walks each day.

her total for the week will be 5*x.